OpenStudy (anonymous):

Rewrite as an equivalent logarithmic expression x= the square root of pr over q^3

OpenStudy (anonymous):

\[\Large {\color{black} {x=\sqrt{q^{3}}}}\]Is that your expression?

OpenStudy (anonymous):

pr is the square root and it is over q cubed

OpenStudy (anonymous):

i can't understand , can you draw it?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\Large x=\frac{\sqrt{pr}}{q^{3}}\] \[\Large x=\frac{(pr)^{1/2}}{q^{3}}\] taking log both sides, \[\Large log x=log[\frac{(pr)^{1/2}}{q^{3}}]\] now apply the property log(m/n) = logm - logn

OpenStudy (anonymous):

Thanks so much you are a life saver.

OpenStudy (anonymous):

:)